OpenStudy (anonymous):

Consider the functions f(x) = 8x + 2 and g(x) = 5x − 3. What is the value of (f/g a.k.a f over g)(4)

OpenStudy (anonymous):

Heres my new question

OpenStudy (anonymous):

please show work

OpenStudy (anonymous):

put a fraction bar between them i.e. write \(\frac{f}{g}(x)=\frac{f(x)}{g(x)}\)

OpenStudy (anonymous):

find \(f(4)\) and find \(g(4)\) and put \(f(4)\) over \(g(4)\)

OpenStudy (anonymous):

f/g = (8x+2)/(5x-3) now find f/g of 4 8(4) + 2/ 5(4) - 3 = 2

OpenStudy (anonymous):

is that the answer fortune_paul?

OpenStudy (anonymous):

yeah thats the answer

OpenStudy (anonymous):

this

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

Lol yeah

OpenStudy (anonymous):

best response

OpenStudy (anonymous):

im posting a new last question. plz answer

OpenStudy (anonymous):

okay..